the length of the rectangle is 3 meters more than its width.If the perimeter is 54 meters,what are its dimensions?

Answer by stanbon(75887)

the length of the rectangle is 3 meters more than its width.If the perimeter is 54 meters,what are its dimensions?
---------------
Let L = W+3
-----
Perimeter = 2(L+W)
Substitute and solve for "W":
54 = 2(W+3 + W)
27 = 2W+3
2W = 24
Width = 12 meters
---
Solve for "L":
Length = W + 3
Length = 15 meters
===========================
Cheers,
Stan H.